OpenStudy (anonymous):

If the point P(-4/5,y) lies on the unit circle and P is in the third quadrant, what does y equal? If necessary, use the slash mark ( / ) for a fraction bar.

OpenStudy (anonymous):

\[-\frac{3}{5}\]

jimthompson5910 (jim_thompson5910):

Use the fact that \[x^2+y^2=1\] is the equation of the unit circle. We know that x=-4/5, so plug this in and solve for y \[x^2+y^2=1\] \[(-\frac{4}{5})^2+y^2=1\] \[\frac{16}{25}+y^2=1\] \[y^2=1-\frac{16}{25}\] \[y^2=\frac{9}{25}\] \[y=-\sqrt{\frac{9}{25}}\] Note: we're in the 3rd quadrant, so y is negative. So we use the negative square root. \[y=-\frac{3}{5}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

jimthompson is correct and so is the method. however, every math teachers favorite right triangle is 3 - 4 - 5. this is a 3, 4 5 right triangle which is why you can do it in your head after you have seen a few

OpenStudy (anonymous):

if you divide 3, 4, 5 by 5 you get \[\frac{3}{5},\frac{4}{5},1\] so it is the same triangle just scaled so that the hypotenuse is 1

OpenStudy (anonymous):

you will see this often along with 5, 12, 13 scaled as \[\frac{5}{13},\frac{12}{13},1\]